### v1

@@ 6.1. Tangent Oseen equations
Regarding the pressure, we claim that there exists a unique $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\overline{\mathbf{u}})\mathbf{u}\cdot\mathbf{v}-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\quad\forall\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma).$ (6.13)
Proceeding as in the proof of 4.5, we notice that because of 4.3 (Helmholtz decomposition), (6.13) is equivalent to
$-\int_{\Gamma}\uppi\,\Delta_{\Gamma}\phi$ $=\langle\mathbf{f},\nabla_{\Gamma}\phi\rangle-\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\nabla_{\Gamma}\phi-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\nabla_{\Gamma}\phi$ (6.14)
$=\langle\mathbf{f},\nabla_{\Gamma}\phi\rangle-\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\nabla_{\Gamma}\phi-\langle K(\mathbf{u}),\nabla_{\Gamma}\phi\rangle=:\langle h,\phi\rangle,\qquad\forall\phi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma).$
Consequently, the existence of a unique $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ follows from the well-posedness of the ultra-weak formulation of the Laplace–Beltrami operator (Section 3) with right-hand side datum $h\in(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}$. Moreover, the following a-priori bound holds true
@@
The combination of this latter estimate with (6.12) results in (6.8).
Finally, if $p=2$ we choose $\mathbf{v}=\mathbf{u}$ in (6.10) and use integration-by-parts formula [11, Lemma 3.3] to obtain
$\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}^{2}=\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}^{2}+\int_{\Gamma}(\nabla_{\Gamma}\overline{\mathbf{u}})\mathbf{u}\cdot\mathbf{u}=\langle\mathbf{f},\mathbf{u}\rangle\leq\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}\|\mathbf{u}\|_{1,\Gamma},$
which combined with 3.8 (Poincaré inequality) for $p=2$, yields the estimate for $\mathbf{u}$ in(6.9). The estimate for $\uppi$ in (6.9) follows from that of $\mathbf{u}$ and (6.15). ∎
@@ 6.2. Existence of solutions of the incompressible tangent Navier–Stokes equations
In this section we establish the existence of solutions to the tangent incompressible Navier–Stokes equations (6.1) (i.e. with $g=0$) for dimensions $d\in\{2,3,4\}$ and $p\geq 2$; see 6.5 below. For the case $p=2$ ($\mathrm{L}^{2}$-based) we utilize the standard Galerkin method based upon eigenfunctions of the Stokes operator (cf. Section 5) to recover a velocity field $\mathbf{u}$. Given the lack of boundary of $\Gamma$, we recover the pressure field $\uppi$ in terms of $\mathbf{u}$ and $\mathbf{f}$ by solving a Laplace–Beltrami problem for $\uppi$. In turn, the case $p>2$ follows from the case $p=2$ in conjunction with the well-posedness of the tangent Oseen equations (cf. 6.2).
@@
Let $d\in\{2,3,4\}$ and assume $\Gamma$ is of class $C^{3}$. Then, for each $\mathbf{f}\in(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ there exists $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ such that
@@
We proceed by a Galerkin method. Let $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}\subseteq\bigcap_{2\leq p<\infty}\mathbf{W}^{2,p}_{t,\sigma}(\Gamma)$ the family of functions provided by 5.3 (eigendecomposition of the Stokes operator). For each fixed $n\in\mathrm{N}$, we seek an approximated velocity $\mathbf{u}_{n}\in\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{n}:\nabla_{\Gamma}\mathbf{v}_{i}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{v}_{i}=\langle\mathbf{f},\mathbf{v}_{i}\rangle,$ (6.17)
for each $i\in\{1,\dotsc,n\}$.
Step 1: Existence of Galerkin solutions. The existence of solutions of (6.17) will follow from the abstract result [75, Page 110, Lemma 1.4], whose hypotheses we now verify. Let $X:=\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$ be the finite dimensional vector space endowed with the inner product $\langle\mathbf{u},\mathbf{v}\rangle_{X}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ (cf. 3.8). Let $P:X\rightarrow X$ be the operator defined by
$\langle P(\mathbf{u}),\mathbf{v}\rangle_{X}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}-\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{u},\mathbf{v}\in X,$
which is clearly continuous. Moreover, by noticing $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{u}=0$ for each $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ (because of integration-by-parts formula [11, Lemma 3.3] and density arguments) and resorting to 3.8 (Poincaré inequality), we have that
$\langle P(\mathbf{u}),\mathbf{u}\rangle_{X}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{u}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{u}-\langle\mathbf{f},\mathbf{u}\rangle=\|\mathbf{u}\|_{X}^{2}-\langle\mathbf{f},\mathbf{u}\rangle\\ \geq\|\mathbf{u}\|_{X}^{2}-\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}\|\mathbf{u}\|_{1,\Gamma}\geq\|\mathbf{u}\|_{X}(\|\mathbf{u}\|_{X}-c\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}),\qquad\forall\mathbf{u}\in X.$
@@
$\langle P(\mathbf{u}),\mathbf{u}\rangle_{X}>0,\qquad\forall\mathbf{u}\in X,\quad\|\mathbf{u}\|_{X}=K.$
This finishes the verification of the hypotheses of [75, Page 110, Lemma 1.4], whence we deduce that there exists $\mathbf{u}\in X$ such that $P(\mathbf{u})=\textbf{0}$, or equivalently, that the Galerkin system (6.17) admits a solution $\mathbf{u}_{n}\in\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$.
Step 2: Passage to the limit. Since $\mathbf{u}_{n}\in\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$ and (6.17) is linear in $\mathbf{v}_{i}$, we can take $\mathbf{v}_{i}\leftarrow\mathbf{u}_{n}$ in (6.17) and use again [11, Lemma 3.3], to arrive at
$\|\nabla_{\Gamma}\mathbf{u}_{n}\|_{0,\Gamma}^{2}=\|\nabla_{\Gamma}\mathbf{u}_{n}\|_{0,\Gamma}^{2}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{u}_{n}=\langle\mathbf{f},\mathbf{u}_{n}\rangle\leq\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}\|\mathbf{u}\|_{1,\Gamma}.$
Hence, by 3.8 (Poincaré inequality), we obtain the a-priori estimate
$\|\mathbf{u}_{n}\|_{1,\Gamma}\leq C\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}},$ (6.18)
where $C>0$ is independent of $n$. Since the sequence of Galerkin solutions $\{\mathbf{u}_{n}\}_{n\in\mathrm{N}}$ is bounded in the Hilbert $\mathbf{H}^{1}_{t,\sigma}$, it follows from the the Banach–Alaoglu Theorem [12, Theorem 3.18] and 2.10 (Sobolev compact embedding) that there exists $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ such that (up to a subsequence)
@@
$\mathbf{u}_{n}$ $\to\mathbf{u},\qquad\text{in $\mathbf{L}^{2}(\Gamma)$}.$ (6.20)
For any $\mathbf{w}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ we write $\mathbf{w}=\sum_{i=1}^{\infty}\alpha_{i}\mathbf{v}_{i}$, with the series converging in the $\mathbf{H}^{1}(\Gamma)$-norm. Denoting $\mathbf{w}_{N}:=\sum_{i=1}^{N}\alpha_{i}\mathbf{v}_{i}\in\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{N}$ for fixed $N\leq n$, we deduce from (6.17) that
@@
Regarding the nonlinear term in (6.21), we first notice that the Sobolev embedding $\mathbb{W}^{1,p}(\Gamma)\hookrightarrow\mathbb{L}^{\infty}(\Gamma)$ for $p>1$ big enough (cf. 2.9) implies that $\nabla_{\Gamma}\mathbf{w}_{N}\in\bigcap_{2\leq p<\infty}\mathbb{W}^{1,p}(\Gamma)\subseteq\mathbb{L}^{\infty}(\Gamma)$. Thus, taking into account the incompressibility condition $\operatorname{div}_{\Gamma}\mathbf{u}_{n}=0$, integration-by-parts formula [11, Lemma 3.3], the strong convergence (6.20) and $\lim_{N\to\infty}\nabla_{\Gamma}\mathbf{w}_{N}=\mathbf{w}$ in $\mathbb{L}^{2}(\Gamma)$ and the fact that $\mathbf{u}\otimes\mathbf{u}\in\mathbb{L}^{2}(\Gamma)$, it transpires that
$\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{w}_{N}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}_{N})\mathbf{u}_{n}\cdot\mathbf{u}_{n}\xrightarrow{n\to\infty}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}_{N})\mathbf{u}\cdot\mathbf{u}\xrightarrow{N\to\infty}-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w})\mathbf{u}\cdot\mathbf{u}=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{w}$ (6.23)
Finally, since $\mathbf{w}_{N}\to\mathbf{w}$ in $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$, it also holds that
@@
Overall, combining (6.21), (6.22), (6.23) and (6.24) yields (6.16) and finishes the proof. ∎
@@
Let $d\in\{2,3,4\}$, $2\leq p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of (6.1) with $g=0$. Moreover, the following a-priori bounds hold for $p=2$:
@@
We first consider the case $p=2$. From 6.4 (existence of a velocity field) we deduce the existence of $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ that satisfies (6.16) with the given $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}\subseteq(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$:
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{H}^{1}_{t}(\Gamma).$ (6.26)
Then, noticing that $\mathbf{u}\in\mathbf{L}^{d}_{t,\sigma}(\Gamma)$ because of 2.6 (Gagliardo–Nirenberg–Sobolev and $d\leq 4$), a direct application of 6.2 (well-posedness of the Oseen equations) with $\overline{\mathbf{u}}$ replaced by $\mathbf{u}$ yields the existence of the pressure $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ such that the pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ solves (6.1) with $g=0$. Moreover, estimates (6.25) follow from the first estimate in (6.9), and the combination of the case $p=2$ of estimate (6.15) with the continuous injection $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{d}(\Gamma)$. ∎
@@
$\|\uppi\|_{0,p;\Gamma}\leq C_{3}\left(\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+(1+\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}})\|\mathbf{u}\|_{1,p;\Gamma}\right),$ (6.27)
@@
We find it important to mention that, in spite of 6.5 not providing a unique pair $(\mathbf{u},\uppi)$, the pressure $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ is uniquely determined by the velocity field $\mathbf{u}$ and right-hand side $\mathbf{f}$. This is due to (6.14) with $\overline{\mathbf{u}}=\mathbf{u}$ within the proof of 6.2 (well-posedness of the Oseen equations).

### v2

@@ 5.1. Tangent Oseen equations
We consider now $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and deal with the pressure field $\uppi$. We claim that there exists a unique $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\quad\forall\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma).$ (5.15)
Proceeding as in the proof of 4.5, we notice that because of 4.3 (Helmholtz–Weyl decomposition on $\Gamma$), (5.15) is equivalent to
$\int_{\Gamma}\uppi\,\Delta_{\Gamma}\phi=\langle\mathbf{f},\nabla_{\Gamma}\phi\rangle-\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\nabla_{\Gamma}\phi-\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\nabla_{\Gamma}\phi\rangle=:\langle h,\phi\rangle,\qquad\forall\phi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma).$ (5.16)
Consequently, the existence of a unique $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ follows from 2.17 (ultra-weak solutions of the Laplace–Beltrami operator on $\Gamma$) with right-hand side datum $h\in(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}$. Moreover, the following a-priori bound holds true
@@
Combining (5.17) with (5.14) results in (5.11).
Finally, if $p=2$ we choose $\mathbf{v}=\mathbf{u}$ in (5.13) and notice that $\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{u}\rangle=0$, to obtain
$\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}^{2}=\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}^{2}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{u}=\langle\mathbf{f},\mathbf{u}\rangle\leq\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}\|\mathbf{u}\|_{1,\Gamma},$
which combined with 3.3 (Poincaré inequality) for $p=2$, yields the estimate for $\mathbf{u}$ in (5.12). The estimate for $\uppi$ in (5.12) follows from that of $\mathbf{u}$ and (5.17). ∎
@@ 5.2. Existence of solutions of the incompressible tangent Navier–Stokes equations
In this section we establish the existence of solutions to the tangent incompressible Navier–Stokes equations (5.1) (i.e. with $g=0$) for dimensions $d\in\{2,3,4\}$ and $p\geq 2$; see 5.5 below. For the case $p=2$ ($\mathrm{L}^{2}$-based) we utilize the standard Faedo–Galerkin method based upon eigenfunctions of the Stokes operator (cf. Section 4.4) to recover a velocity field $\mathbf{u}$. Given the lack of boundary of $\Gamma$, we recover the pressure field $\uppi$ in terms of $\mathbf{u}$ and $\mathbf{f}$ by solving a Laplace–Beltrami problem for $\uppi$. Moreover, the case $p>2$ follows from the case $p=2$ in conjunction with the well-posedness of the tangent Oseen equations (cf. 5.2).
@@
Let $d\in\{2,3,4\}$ and assume $\Gamma$ is of class $C^{2}$. Then, for each $\mathbf{f}\in(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ there exists $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ such that
@@
We proceed by a Faedo–Galerkin method. Let $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}\subseteq\bigcap_{2\leq p<\infty}\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ be the family of functions provided by 4.10 (eigendecomposition of the Stokes operator). For each fixed $n\in\mathrm{N}$, we seek an approximated velocity $\mathbf{u}_{n}\in\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{n}:\nabla_{\Gamma}\mathbf{v}_{i}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{v}_{i}=\langle\mathbf{f},\mathbf{v}_{i}\rangle,\qquad\forall i\in\{1,\dotsc,n\}.$ (5.19)
Step 1: Existence of Galerkin solutions. The existence of solutions of (5.19) will follow from the abstract result [63, p. 110, Lemma 1.4], whose hypotheses we now verify. Let $X_{n}:=\operatorname{span}\{\mathbf{v}_{i}\}_{i=1}^{n}$ be the finite dimensional vector space endowed with the inner product $(\mathbf{u},\mathbf{v})_{H}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}$ from (4.19). Let $P:X_{n}\rightarrow X_{n}$ be the operator defined by
$(P(\mathbf{u}),\mathbf{v}\rangle_{H}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}-\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{u},\mathbf{v}\in X_{n},$
which is clearly continuous. Moreover, by noticing $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{u}=0$ for each $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ (because of identity (5.6) and density arguments) and resorting to 3.3 (Poincaré inequality), we have that
$(P(\mathbf{u}),\mathbf{u})_{H}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{u}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{u}-\langle\mathbf{f},\mathbf{u}\rangle=\|\mathbf{u}\|_{H}^{2}-\langle\mathbf{f},\mathbf{u}\rangle\\ \geq\|\mathbf{u}\|_{H}^{2}-\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}\|\mathbf{u}\|_{1,\Gamma}\geq\|\mathbf{u}\|_{H}(\|\mathbf{u}\|_{H}-c\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}),\qquad\forall\mathbf{u}\in X_{n}.$
@@
$(P(\mathbf{u}),\mathbf{u})_{H}>0,\qquad\forall\mathbf{u}\in X_{n},\quad\|\mathbf{u}\|_{H}=K.$
This finishes the verification of the hypotheses of [63, p. 110, Lemma 1.4], whence we deduce that there exists $\mathbf{u}\in X_{n}$ such that $P(\mathbf{u})=\textbf{0}$, or equivalently, that the Galerkin system (5.19) admits a solution $\mathbf{u}_{n}\in X_{n}$.
Step 2: Passage to the limit. Since $\mathbf{u}_{n}\in X_{n}$ and (5.19) is linear in $\mathbf{v}_{i}$, we can take $\mathbf{v}_{i}\leftarrow\mathbf{u}_{n}$ in (5.19) and use again (5.6), to arrive at
$\|\nabla_{\Gamma}\mathbf{u}_{n}\|_{0,\Gamma}^{2}=\|\nabla_{\Gamma}\mathbf{u}_{n}\|_{0,\Gamma}^{2}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{u}_{n}=\langle\mathbf{f},\mathbf{u}_{n}\rangle\leq\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}\|\mathbf{u}_{n}\|_{1,\Gamma}.$
Hence, by 3.3 (Poincaré inequality), we obtain the a-priori estimate
$\|\mathbf{u}_{n}\|_{1,\Gamma}\leq C\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}},\qquad\forall n\in\mathrm{N},$
where $C>0$ is independent of $n$. Since the sequence of Galerkin solutions $\{\mathbf{u}_{n}\}_{n\in\mathrm{N}}$ is bounded in the Hilbert space $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$, it follows from the the Banach–Alaoglu Theorem [12, Theorem 3.18] and 2.10 (compact Sobolev embedding) that there exists $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ such that (up to a subsequence)
@@
$\mathbf{u}_{n}$ $\to\mathbf{u},\qquad\text{in $\mathbf{L}^{4-\varepsilon}(\Gamma)$},$ (5.21)
for some arbitrary, but fixed, $\varepsilon\in(0,1)$. Notice that $\varepsilon$ can be taken as 0 if $d\in\{2,3\}$. On the other hand, even for $d=4$, $\mathbf{u}\in\mathbf{L}^{4}(\Gamma)$ because of the continuous injection $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{4}(\Gamma)$.
For any $\mathbf{w}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ we write $\mathbf{w}=\sum_{i=1}^{\infty}\alpha_{i}\mathbf{v}_{i}$, with the series converging in the $\mathbf{H}^{1}(\Gamma)$-norm. Denoting $\mathbf{w}_{N}:=\sum_{i=1}^{N}\alpha_{i}\mathbf{v}_{i}\in X_{N}$ for fixed $N\leq n$, we deduce from (5.19) that
@@
Regarding the nonlinear term in (5.22), we first notice that using the strong convergence (5.21) it is easy to see that that $\mathbf{u}_{n}\otimes\mathbf{u}_{n}\to\mathbf{u}\otimes\mathbf{u}$ in $\mathbb{L}^{\frac{4-\varepsilon}{2}}(\Gamma)$. Thus, taking also into account the incompressibility condition $\operatorname{div}_{\Gamma}\mathbf{u}_{n}=0$ and identity (5.6) with $\overline{\mathbf{u}}=\mathbf{u}_{n}$ and the fact that $\nabla_{\Gamma}\mathbf{w}_{N}\in\mathbb{L}^{p}(\Gamma)$ for every $p\in(1,\infty)$, it first transpires that
$\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u}_{n})\mathbf{u}_{n}\cdot\mathbf{w}_{N}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}_{N}):\mathbf{u}_{n}\otimes\mathbf{u}_{n}\xrightarrow{n\to\infty}-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}_{N}):\mathbf{u}\otimes\mathbf{u}.$ (5.24)
Moreover, because $\mathbf{u}\otimes\mathbf{u}\in\mathbb{L}^{2}(\Gamma)$ and $\nabla_{\Gamma}\mathbf{w}_{N}\to\nabla_{\Gamma}\mathbf{w}$ in $\mathbb{L}^{2}(\Gamma)$ as $N\to\infty$, identity (5.6) with $\overline{\mathbf{u}}=\mathbf{u}$, yields
$\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}_{N}):\mathbf{u}\otimes\mathbf{u}\xrightarrow{N\to\infty}-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{w}):\mathbf{u}\otimes\mathbf{u}=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{w}.$ (5.25)
Finally, since $\mathbf{w}_{N}\to\mathbf{w}$ in $\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ as $N\to\infty$, it also holds that
@@
Overall, combining (5.22), (5.23), (5.24), (5.25) and (5.26) yields (5.18) and finishes the proof. ∎
@@
Let $p\in[2,\infty)$ and assume $\Gamma$ is of class $C^{2}$ and of dimension $d\in\{2,3,4\}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of (5.1) with $g=0$. Moreover, the following a-priori bounds hold for $p=2$:
@@
From 5.4 (existence of a velocity field) we deduce the existence of $\mathbf{u}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ that satisfies (5.18) with the given $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\subseteq(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$:
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{H}^{1}_{t}(\Gamma).$
Let $r=r(\varepsilon,p,d)$ be given by (5.7) for some fixed $\varepsilon>0$. Then, noticing that $\mathbf{u}\in\mathbf{L}^{r}_{t,\sigma}(\Gamma)$, because of 2.8 (Gagliardo–Nirenberg–Sobolev) and $d\leq 4$, a direct application of 5.2 (well-posedness of the Oseen equations) with $\overline{\mathbf{u}}$ replaced by $\mathbf{u}$ yields the existence of the pressure $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$, and the higher regularity $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ and the fact that the pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ solves (5.1) with $g=0$. Moreover, estimates (5.27) follow from the first estimate in (5.12), and combining the case $p=2$ of estimate (5.17) with the continuous injection $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{d}(\Gamma)$. ∎
@@
$\|\uppi\|_{0,p;\Gamma}\leq C_{3}\left(\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+(1+\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}})\|\mathbf{u}\|_{1,p;\Gamma}\right),$
@@
We find it important to mention that, in spite of 5.5 not providing a unique pair $(\mathbf{u},\uppi)$, the pressure $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ is uniquely determined by the velocity field $\mathbf{u}$ and right-hand side $\mathbf{f}$. This is due to (5.16) with $\overline{\mathbf{u}}=\mathbf{u}$ within the proof of 5.2 (well-posedness of the Oseen equations).